# ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models

### v1

@@ 5.4 Ablation Studies
5.4.1 Using Smaller Budgets
In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls, mainly following previous work Lin et al. (2024); Chen et al. (2024), avoiding any potential advantage that could arise from optimizing this number. In the Appendix, we provide reward plots for the ACING approach, showing the best achieved reward over the calls. As shown in various plots in Figure 3, the ACING approach found the optimal instruction (achieving a reward of 1) within just a few black-box calls. Some tasks required fewer than 10 API calls to find the optimal instruction, as in the case of ‘active to passive’ and ‘letters list’, and fewer than 20 calls for tasks like ‘translation’ and ‘diff’. It can be seen that the vast majority of tasks achieved their best reward value within the first 60 to 80 calls, demonstrating that ACING can even be used for more constrained budgets.
5.4.2 Using a Different White-box Model
In the main paper, we present results using Vicuna as the white-box language model, following the setup of prior work. To evaluate the impact of different white-box models on the performance of our method, we also tested ACING with WizardLM-13B-v1.2 Xu et al. (2024) compared to Vicuna-13B-v1.3. The experiments in Table 8 indicate that WizardLM achieved a higher median test score across all tasks and excelled in a greater number of top-performing tasks, demonstrating that tuning the choice of the white-box model can improve results.
5.4.3 Using Different Action Dimensionalities
In the main paper, we present results using actions with dimension $d^{\prime}=10$, following the setup of prior work. To evaluate the performance of ACING with different dimensionalities, we conducted an experiment using $d^{\prime}\in\{5,10,20,40,100\}$, keeping other parameters fixed, for a budget of $165$ calls. The results, shown in Table 6, indicate that while the smallest dimension, $d^{\prime}=5$, recovered the best scores for some tasks, it generally has the lowest performance across most tasks. Furthermore, both $d^{\prime}=10$ and $d^{\prime}=20$ yield similar performance in terms of the number of best-performing tasks, indicating a low sensitivity to this parameter. For the larger dimension, $d^{\prime}=40$, the method achieved the highest number of best-performing tasks, showing that tuning (increasing) $d^{\prime}$ can improve results.
5.4.4 Using a Different Number of Exemplars
In the main paper, we present results using five exemplars, following the setup of prior work. To evaluate the performance of ACING with a smaller number of exemplars, we conducted an experiment using a single exemplar, i.e., $|\mathcal{E}|=1$. Results shown in Table 7 confirm that using five exemplars generally yield higher test scores and a greater number of best-performing tasks than a single exemplar. However, a single exemplar still performs well, matching the performance of five exemplars across several tasks (e.g., phonetics, summation, morpho-syntax, translation) and sometimes even outperforming them. Nonetheless, five exemplars offer an advantage, particularly for more cognitively complex tasks.
@@ 6 Related Work
Many early approaches such as AutoPrompt Shin et al. (2020), FluentPrompt Shi et al. (2023), and other soft prompt-based methods Lester et al. (2021); Li and Liang (2021); Zhong et al. (2021) focused on optimizing prompts for white-box LLMs and cannot be directly applied to black-box LLMs. Moreover, several techniques have also been proposed for optimizing prompts for grey-box LLMs: BBT Sun et al. (2022b), BBTv2 Sun et al. (2022a), and clip-tuning Chai et al. (2022) use evolutionary algorithms for this purpose. However, these methods still require access to input token embeddings and output logits, making them unsuitable for settings where only query access is allowed.
Other approaches Li and Liang (2021); Zhong et al. (2021) employ continuous soft prompts by fine-tuning parameters of specific input tokens, but these often generate human-unreadable prompts and are impractical for API-accessible LLMs. Some other methods Deng et al. (2022); Zhang et al. (2023), created for optimizing discrete prompts instead of continuous ones, can optimize discrete tokens but require access to the LLM’s output distribution, which is typically unavailable in black-box LLMs. Zhou et al. (2023) proposed generating candidate prompts through resampling without specific optimizations, while Guo et al. (2024) extended this model-free approach by using evolutionary algorithms. However, these methods usually require a large number of queries.
Instruction optimization in black-box LLMs recently focused on searching through the space of soft prompts, then relying on open-source white-box LLMs as a means to transform these soft prompts into discrete prompts. InstructZero Chen et al. (2024) learns discrete prompts for ChatGPT by searching through the continuous space of soft prompts using Bayesian optimization, which are then transformed into discrete prompts by a white-box LLM. INSTINCT Lin et al. (2024), a successor to this approach, relies on the same approach. However, instead of Bayesian optimization, it uses NeuralUCB Zhou et al. (2020), a contextual bandit (selection) algorithm that searches within a finite set of continuous actions generated using Sobol Eriksson et al. (2019). Our approach combines the hybrid strategy of using white-box and black-box LLMs with an actor-critic approach that operates in a stateless, infinite continuous-action space.
TRIPLE Shi et al. (2024) modeled the prompt learning problem as a best-arm-identification problem. However, it only considered the problem as a selection task, assuming a given pool of discrete prompts. In contrast, our method addresses both the generation and selection of prompts.
@@ 7 Conclusion
ACING presents a promising approach for learning prompts for black-box LLMs by framing it as a stateless continuous-action RL problem. Our method shows significant improvements over recent baseline methods. Evaluations across 31 datasets reveal that ACING enhances prompt quality, achieving a median score improvement of 10 percentage points and outperforming human-crafted instructions by up to 39 percentage points. This sets a new benchmark for automated instruction optimization in black-box LLMs.
@@ 8 Limitations
While our proposed results were achieved using default and commonly used values for hyperparameters in network design, learning rates, and following previous works for some common parameters, the results presented depend on these choices. These results could be further improved with fine-tuning, as demonstrated in the ablation studies (§ 5.4). Another limitation of this work, similar to previous studies, is the risk of overfitting when dealing with a limited number of validation examples. This can result in a prompt that minimizes the loss on the validation set but does not generalize well to the test dataset, leading to a discrepancy between training and test performance. Another limitation of this work, similar to previous work (Lin et al., 2024; Chen et al., 2024), is the reliance on a white-box model, where the choice of model may lead to performance variations, as demonstrated in § 5.4.2. Avoiding this limitation requires acting directly on the black-box LLM, which involves handling actions in a large discrete action space. This necessitates adapting current RL approaches for large discrete action spaces, as studied in Fourati et al. (2024) and the references therein.
Appendix A Experimental Details
A.1 Hyperparameters Remark
Across the diverse tasks, in the main paper, the same hyperparameters were used, which shows that the algorithm generalizes well across the 30 tasks without specifically tuning hyperparameters in each task. A summary of the key parameters cab be found in the following Table. An implementation of ACING is available at https://github.com/salmakh1/ACING.
@@ A.1 Hyperparameters Remark
White-box $h$ Vicuna-13B
Actor-network $(1-1024-256-10)$
Critic-network $(10-128-128-1)$
@@
Intrinsic (action) dimension $d^{\prime}$ 5
@@
A.2 Actor-Critic Details

### v2

@@ 4.4 Ablation Studies
We present detailed ablation studies on key design choices, summarize some below, and provide full results and plots in Appendices F.1F.7.
Use of critics. We compare our method (with two critics) to variants with a single critic and to a baseline without a critic (policy-gradient). As shown in Appendix F.1, the two-critic architecture yields the highest accuracy, best stability, and top performance in difficult settings (e.g., cognitive), confirming the value of conservative estimation.
Budget efficiency. While ACING uses a fixed 165-query budget for fair comparison with prior work, it often converges well before the budget is exhausted. As illustrated in Appendix F.3, many tasks reach optimal rewards within 60–80 queries, and some within 10–20, showing strong sample efficiency under constrained settings.
White-box model. Using WizardLM-13B Xu et al. (2024) instead of Vicuna improves median test accuracy by 8 points and increases the number of best-performing tasks (Appendix F.7). This demonstrates that ACING can benefit from stronger decoding models, although it remains effective across architectures.
Action dimensionality. We test latent action sizes $d^{\prime}\in\{5,10,20,40,100\}$. Results in Appendix F.5 show that $d^{\prime}=10$ and $d^{\prime}=20$ perform consistently well, while larger dimensions like $d^{\prime}=40$ yield improvements on specific tasks.
Number of exemplars. Using a single exemplar performs surprisingly well and matches the 5-exemplar setup on several tasks (e.g., phonetics and summation). Still, five exemplars offer more consistent gains on complex tasks (Section˜F.6).
Budget splitting. We explore dividing the query budget into an exploration phase and a final re-ranking phase, where top prompts are re-evaluated multiple times. This two-phase strategy improves median test scores by 5 points and boosts the number of best-performing tasks (Section˜F.4).
Figure 3: Ranking distributions across cognitive tasks for all algorithms, ordered increasing by median rank.
@@ 5 Related Work
Early approaches such as AutoPrompt Shin et al. (2020), FluentPrompt Shi et al. (2023), and soft prompt tuning Lester et al. (2021); Li and Liang (2021); Zhong et al. (2021) rely on access to model gradients or embeddings, limiting their applicability in black-box settings. Grey-box methods such as BBT and BBTv2 Sun et al. (2022b, a) and CLIP-tuning Chai et al. (2022) relax these constraints by leveraging token embeddings or logits, but are still incompatible with API-only black-box models.
Several recent works frame instruction search as a discrete optimization problem. RLPrompt Deng et al. (2022) and Tempera Zhang et al. (2023) use RL to identify prompts, but assume access to token-level outputs or confidence scores. Alternatively, sampling- and evolution-based strategies such as APE Zhou et al. (2023), PromptBreeder Fernando et al. (2024), EvoPrompt Guo et al. (2024), PromptWizard Agarwal et al. (2024), and Auto Evol-Instruct Zeng et al. (2024) iteratively generate and refine candidate prompts via LLM sampling or mutation. While effective in some cases, these methods typically rely on large candidate pools or expensive query budgets.
Zeroth-order optimization (e.g., AIO Qi et al. (2025), ZOPO Hu et al. (2024)) estimates gradients without backpropagation, yet incurs high computational cost and query complexity. InstructZero Chen et al. (2024) and INSTINCT Lin et al. (2024b) apply Bayesian Optimization (BO) and NeuralUCB Zhou et al. (2020) within a finite action space. In contrast, StablePrompt Kwon et al. (2024) uses PPO to stabilize discrete prompt learning, but fine-tune white-box LLMs. Our approach differs by formulating prompt optimization as a continuous-action RL problem, enabling more efficient exploration without gradient access or large models.
Complementary strategies include exemplar selection Wu et al. (2024), preference-based feedback Lin et al. (2024a), and best-arm identification Shi et al. (2024). However, these typically rely on fixed prompt pools or require human preference labels. ACING jointly addresses both the generation and selection in a unified framework, yielding high-performing prompts without pool constraints or external supervision.
@@ 6 Conclusion
We present ACING, an actor-critic RL framework for prompt optimization in black-box LLMs. By formulating the task as a continuous-action problem, ACING enables exploration of an infinite space through an entropy-regularized policy under strict query constraints. It outperforms strong baselines and surpasses human-written instructions, without any per-task tuning, demonstrating effectiveness, efficiency, and practicality.
@@ 7 Limitations
While ACING achieves strong and consistent performance across diverse tasks, we do not claim universal superiority on every individual task. Given the stochastic and non-convex nature of prompt optimization—particularly in black-box LLMs under tight query budgets—strong prompts can occasionally arise even from random search or other optimization approaches. However, ACING substantially improves the chances of discovering such prompts through a principled exploration strategy grounded in actor-critic learning.
Similar to previous prior works (Lin et al., 2024b; Chen et al., 2024; Hu et al., 2024), our reliance on a white-box model introduces variability, as model selection affects performance (Appendix F.7). Addressing this limitation would require operating directly on black-box LLMs, which poses challenges due to the large discrete action space. Adapting RL methods for such spaces, as explored in Fourati et al. (2024), and proposing hybrid approaches combining soft-prompt and discrete optimization, represent promising directions for future research.
Despite the strong performance achieved with a fixed hyperparameter configuration across all tasks, we acknowledge that domain-specific tuning might yield marginal gains, particularly in edge cases. That said, our goal was to emphasize generality and simplicity without introducing per-task overhead—an important factor for real-world usability.
8 Ethical Considerations
ACING automates prompt optimization for LLMs, which can reduce manual effort but may also amplify risks associated with bias, misuse, or harmful content generation. We emphasize that ACING is a general optimization framework and does not guarantee ethical outputs from the underlying LLMs. It should be paired with appropriate content filters and safety mechanisms when deployed in sensitive domains. Optimized prompts could reinforce or amplify existing biases present in the base LLMs. Future work should incorporate fairness-aware reward functions or post-hoc bias mitigation.
Lastly, our experiments assume legal and ethical API usage, and we caution against applying instruction optimization to restricted or proprietary models without adherence to usage policies and terms of service.
Appendix A Table of Notations
For clarity, we summarize all symbols used in our actor-critic framework in Table 6.
Symbol Meaning
$f$ Black-box LLM (only accessible via API)
$h$ White-box LLM (frozen, used as decoder)
$D$ Task distribution over input–output pairs $(x,y)$
$V$ Validation set $\{(x_{j},y_{j})\}_{j=1}^{m}$
$T$ Test set $\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$
$q(\cdot,\cdot)$ Scoring function measuring output quality
$\tau$ Discrete prompt
$\tau(z)$ Prompt generated from soft prompt $z$
$\tau(Pa)$ Prompt from projected action vector $a$
$\tau^{\star}$ Optimal prompt
$z\in\mathbb{R}^{d}$ Soft prompt vector (continuous)
$a\in[0,1]^{d^{\prime}}$ Action (latent prompt representation)
$P\in\mathbb{R}^{d\times d^{\prime}}$ Random projection matrix
$\pi(\cdot;\theta)$ Policy network (actor) with parameters $\theta$
$Q_{w_{1}}(a),Q_{w_{2}}(a)$ Twin critic networks with parameters $w_{1},w_{2}$
$Q(a)$ Value estimate: $Q(a)=\min\{Q_{w_{1}}(a),Q_{w_{2}}(a)\}$
$J_{Q}(w)$ Critic loss: $\mathbb{E}_{a\sim D}\left[\tfrac{1}{2}(Q_{w}(a)-r(a))^{2}\right]$
$J_{\pi}(\theta)$ Actor loss (entropy-regularized policy objective)
$r(a)$ Reward obtained from the environment
$\alpha$ Temperature coefficient for entropy regularization
$H_{\text{target}}$ Target entropy for adaptive $\alpha$
Table 6: Summary of notations used in the ACING actor-critic framework.
Appendix B Experimental Details
B.1 Hyperparameters
Across the diverse tasks, in the main paper, the same hyperparameters were used, which shows that the algorithm generalizes well across the 30 tasks without specifically tuning hyperparameters in each task. A summary of the key parameters can be found in the following Table.
@@ B.1 Hyperparameters
White-box $h$ Vicuna-13B and WizardLM
Actor-network (1102425610)(1-1024-256-10)
Critic-network (101281281)(10-128-128-1)
@@
Intrinsic (action) dimension $d^{\prime}$ 10
@@
Furthermore, like previous works, we use their default (tuned) hyperparameters for the results in the main paper, including the intrinsic dimension $d^{\prime}=10$ and the number of soft tokens $N_{z}=5$. For fairness, we refrain from fine-tuning these parameters for our method and use the same values as in prior works. This ensures that our ACING algorithm searches in the same space, $0,1]^{10}$, and uses the same total number of queries to the black-box LLM as APE, InstructZero, and INSTINCT for a fair comparison. For each algorithm, after identifying the best instruction using the validation set $\mathcal{V}$, we evaluate the discovered instruction on a separate test set $\mathcal{T}$ and report the test score.
B.2 Actor-Critic Details